OpenStudy (anonymous):

calculate angular velocity if its displacement is 90 degree and it complete this displacement in 5 seconds?

OpenStudy (anonymous):

Angular velocity is equal to the displacement over the time it takes to do that displacement. So the formula is \[\omega=\frac{\Delta \theta}{\Delta t} \]. Also you should convert the displacement to radians. Angular velocity is analogous to velocity v is the way that: \[\ v =\frac{\Delta S}{\Delta t} \]Where S - distance traveled; t - time taken See the similarities?

OpenStudy (anonymous):

using knowledge gained by above question convert 90 degrees to radians and rest is same as gogind says then u get answer in radians/second